OpenStudy (anonymous):

expand this binomial (x+3)^5

OpenStudy (amoodarya):

x^5+5(x^4*3)+10(x^3*3^2)+10(x^2*3^3)+5(x*3^4)+3^5 x^5+15x^4+90x^3+270x^2+405x+243

OpenStudy (anonymous):

Use the binomial expansion: (x+3)^5= (3^5)*(1+x)^5 Find the binomial expansion of (1+x)^5, then multiply the answer by 3^5

OpenStudy (anonymous):

i stilll dont know how

OpenStudy (anonymous):

The binomial expansion is given by: \[(1+x)^p=1+px+\frac{(p)(p-1)x^2}{2!} +\frac{(p)(p-1)(p-2)x^3}{3!}+..+...\]

OpenStudy (amoodarya):

do you know the pascal triangle ?

OpenStudy (anonymous):

Btw, mistake on my above comment ... (x+3)^5= (3^5)*(1+x/3)^5

OpenStudy (anonymous):

So use the binomial expansion, and wherever you see 'x', replace that with (x/3)